OpenStudy (anonymous):

Radical equation frac{8}{x(x-2} + frac{2}{x}= frac{4x}{x-2}

OpenStudy (anonymous):

\[ \frac {8}{x(x-2} + \frac{2}{x}= \frac{4x}{x-2}\]

hero (hero):

Multiply both sides by \[x^2 -2x\]

hero (hero):

That will get rid of the denominator

hero (hero):

\[8x^2 - 16x + 2x^2 - 4x = 4x^3-8x^2\]

hero (hero):

\[4x^3-18x^2+20x = 0\]\[2x(2x^2-9x+10)=0\]

OpenStudy (anonymous):

Is that how to check it?

hero (hero):

No, we're finding x

OpenStudy (anonymous):

oh. oops.

hero (hero):

What did you do?

OpenStudy (anonymous):

I'm just trying to follow what you're saying.

hero (hero):

x = 0 is one of the solutions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

How do I find the second solution

hero (hero):

\[2x^2-4x-5x +10 = 0\]\[2x(x-2)-5(x-2) = 0\]\[(x-2)(2x-5)=0\]

hero (hero):

\[x = 2, x = \frac{5}{2}\]

hero (hero):

If you really want to know how I solved all of that you can come to twiddla

hero (hero):

http://www.twiddla.com/solved

OpenStudy (anonymous):

I'll have to check it out soon. thanks

hero (hero):

Well, I won't be around long

hero (hero):

It's a live classroom and I have my own domain. So I can't just post the answer unless you're there to witness. Because other students will visit that domain later on, so I'll have to erase it soon. But if you're not going to be there right now, there's no point in posting it.